OpenStudy (anonymous):

X: 3(x+1) = -2(x+1) -4

OpenStudy (anonymous):

3(x+1) = -2(x+1) -4 3(x+1)+2(x+1) = -4 3x +3 +2x +2 = -4 now combine the like terms and take all the constants on other side of this equation